# Rational solutions of Painlevé V from Hankel determinants and the asymptotics of their pole locations

### v2

@@ 5.1.1 Conclusion of the steepest descent analysis
${\bf G}(z):=\left\{\begin{array}[]{cc}M_{{}_{Q}}(z)&z\not\in\mathbb{D}_{+}\cup\mathbb{D}_{-}\\ M_{{}_{Q}}(z)Q(z)^{\frac{\sigma_{3}}{2}}\begin{bmatrix}1&-1\\ i&i\\ \end{bmatrix}\frac{\displaystyle\zeta_{\pm}(z)^{-\sigma_{3}/4}}{4\sqrt{\pi}}\mathcal{P}_{\pm}(z)&z\in\mathbb{D}_{\pm}\end{array}\right..$ (5.33)
The matrix ${\bf G}(z)$ has the same jumps as $T_{0}(z)$ within $\mathbb{D}_{\pm}$ and on $\Gamma_{m}$; it has additional jump discontinuities across the boundaries $\partial\mathbb{D}_{\pm}$ of the form
${\bf G}(z_{+})={\bf G}(z_{-})\bigg{(}\mathbf{1}+\mathcal{O}(n^{-1})\bigg{)},\ \ \ z\in\partial\mathbb{D}_{\pm}.$ (5.34)
Consequently the matrix
@@
has only jump discontinuities on $\partial\mathbb{D}_{+}\cup\partial\mathbb{D}_{-}\cup\mathscr{L}_{+}\cup\mathscr{L}_{-}$ with all jump matrices that can be verified to be of the form $\mathbf{1}+\mathcal{O}(n^{-1})$, see Fig. 7.
It is then a standard result referred to in the literature as a “small norm theorem” [25, 26] that the solution of this last RHP $\mathscr{E}(z)$ exists and is uniformly close to the identity matrix. Reversing the order of the various transformations, this proves that $\forall s\in{\mathbb{C}}\setminus EoT$, the solution $Y_{n}(z;s)$ of RHP 5.1 exists (at least for $n$ large enough).
More precisely we can state as a result of the above discussion the following:
Theorem 5.2
For any closed set $\mathcal{K}$ in the complement of the closure of $EoT$, there is $N_{0}\in{\mathbb{N}}$ such that the solution of RHP 5.1, and hence no zeroes of $\mathcal{T}_{n}(s)$ (3.1) lie in $\mathcal{K}$ for $n\geq N_{0}$.
We only give a sketch of the proof, which is a relatively standard argument. We summarize the main steps here for the reader with some experience in the Deift-Zhou type of analysis.
As discussed previously in this section, the solvability of RHP 5.1 relies on the solvability of the RHP for the error matrix $\mathscr{E}(z)$ 5.35, whose jump matrices, $J_{\mathscr{E}}(z)$, can be bound uniformly, along the contours of the discontinuities, by
@@
The constant $C$ depends on the radii of the local disks $\mathbb{D}_{\pm}$, which is only limited by the distance of the branch points ±$\pm\frac{is}{2}$ from $z=1$, and it is bounded away from zero uniformly as $s\in\mathcal{K}$. The other source of potential loss of control over the error term is the distance of the main and complementary arcs $\Gamma_{m},\Gamma_{c}$ from $z=1$; indeed if either arc contains $z=1$, the asymptotic analysis would require the construction of a special local parametrix and the small norm theorem cannot be used in the same way. However, as long as $s\in\mathcal{K}$ also the distance of $\Gamma_{m}\cup\Gamma_{c}$ from $z=1$ is bounded away from zero, and hence we can open the lenses as normal. Along the rims of the lenses, $\mathscr{L}_{\pm}$, and along $\Gamma_{c}$, the jump matrices of $\mathscr{E}$ tend exponentially to the identity, so that the main source of error is coming from the boundaries of $\mathbb{D}_{\pm}$.
Thus, in conclusion, as long as both branch points ±$\pm\frac{is}{2}$ and $\Gamma_{c}\cup\Gamma_{m}$ remains at finite positive distance from $z=1$ the constant $C$ in (5.36) can be chosen uniformly with respect to $s\in\mathcal{K}$ so that for $n$ sufficiently large the small norm theorem comes into effect, and uniformly so with respect to $s$.
@@
5.2 Asymptotic analysis inside $EoT$
Figure 8: The contours $\Gamma_{m}$ (red) and $\Gamma_{c}$ (green) for two antipodal values $s_{1}=-s_{2}$ of $s$ with $\mathrm{Re}\,(s_{1})>0$ (left) and $\mathrm{Re}\,(s_{2})<0$ (right). Specifically here $s_{1}=0.48+0.48i=-s_{2}$. The dashed line indicates the branch cut of the function $Q(z)$, which runs partially along $\Gamma_{b}^{1}$. The components of $\Gamma_{c}$ (green arcs) can be chosen freely within the shaded region, which indicates the region where $\mathrm{Re}\,\varphi<0$.
With reference to Fig. 8 we have already observed that the level sets of $\mathrm{Re}\,\varphi(z;s)$ depend only on $s^{2}$ and hence are invariant under the map $s\mapsto-s$; the same is not true for the branch cuts $\Gamma_{m}$ and the regions of positivity of $\mathrm{Re}\,\varphi$. The decision of where to place the branch cuts is forced by the “sign distribution” (i.e. in which regions we have $\mathrm{Re}\,\varphi>0$): this is mandated by the behaviour near infinity $\mathrm{Re}\,\varphi=2\ln|z|+\mathcal{O}(1)$ for the outer region, and by the behaviour near the origin $\mathrm{Re}\,\varphi=\mathrm{Re}\,\frac{s}{z}+\mathcal{O}(1)$ for the two bounded regions. Before we fork the analysis according to the sign of $\mathrm{Re}\,(s)$ we make some common preparation.
As shown in Fig. 8 one of the arcs of $\Gamma_{m}$ consists of an arc joining $z=1,b$, denoted $\Gamma_{b}^{1}$ (the orientation of the arc is suggested by the index placements). Thus, we need to preemptively modify the initial RHP 2.1 so that the contour $\gamma$ is “retracted” in part along $\Gamma_{b}^{1}$. We leave the details to the reader and we report the resulting RHP, which is now the starting point of the following analysis.
Problem 5.5
Let $\gamma=\Gamma_{b}^{1}\cup\gamma_{b}$ consist of the arc $\Gamma_{b}^{1}$ and a loop, $\gamma_{b}$, starting and ending at $z=b$ and containing the origin. Let $Q(z)$ in (5.3) be defined with the branch cut extending from $z=0$ to $z=b$ and then along $\Gamma_{b}^{1}$, as shown in Fig. 8 by the dashed line. Given $n\in\mathbb{N}$, find a $2\times 2$ matrix–valued function $Y(z)=Y_{n}(z)$ such that
  1. 1.
@@ 5.2 Asymptotic analysis inside $EoT$
    $Y(z_{+})$ $=Y(z_{-})\begin{bmatrix}1&Q(z){\rm e}^{\frac{ns}{z}}\\ 0&1\end{bmatrix},\qquad$ $\forall z\in\gamma_{b}$ (5.37)
    $Y(z_{+})$ $=Y(z_{-})\begin{bmatrix}1&\varkappa Q(z_{+}){\rm e}^{\frac{ns}{z}}\\ 0&1\end{bmatrix},\qquad$ $\forall z\in\Gamma_{b}^{1},$ (5.38)
@@
    $\varkappa:={\rm e}^{-2i\pi{{\rho}}}-1.$ (5.39)
  3. 3.
    As $z\to\infty$ the matrix $Y_{n}(z)$ admits an asymptotic expansion of the form (2.18).
The first transformation is the same as in (5.5) and leads to the RHP for $W$ as follows
Problem 5.6
$W(z_{+})=W(z_{-})J_{W}(z),\ \ \ \ z\in\gamma$ (5.40)
$W(z)=\mathbf{1}+\mathcal{O}\left(\frac{1}{z}\right),\ \ \ \ z\to\infty$ (5.41)
@@
$J_{W}(z)$ $=\begin{bmatrix}{\rm e}^{-\frac{n}{2}\left(\varphi(z_{+})-\varphi(z_{-})\right)}&Q(z){\rm e}^{\frac{n}{2}\left(\varphi(z_{+})+\varphi(z_{-})\right)}\\ 0&{\rm e}^{\frac{n}{2}\left(\varphi(z_{+})-\varphi(z_{-})\right)}\end{bmatrix},$ $z\in\Gamma\setminus\Gamma_{b}^{1}$ (5.42)
$J_{W}(z)$ $=\begin{bmatrix}{\rm e}^{-\frac{n}{2}\left(\varphi(z_{+})-\varphi(z_{-})\right)}&\varkappa Q(z_{+}){\rm e}^{\frac{n}{2}\left(\varphi(z_{+})+\varphi(z_{-})\right)}\\ 0&{\rm e}^{\frac{n}{2}\left(\varphi(z_{+})-\varphi(z_{-})\right)}\end{bmatrix},$ $z\in\Gamma_{b}^{1}.$ (5.43)
The second transformation is also similar to (5.9). However, from this point onwards, the details of the transformation depend on the case $\mathrm{Re}\,(s)>0$ or $\mathrm{Re}\,(s)<0$ and thus are given in Sections 5.2.1 and 5.2.2, respectively.
We will not carry out completely the error analysis, which would require the construction of appropriate local parametrices near the points $z=b,a_{\pm},1$. These are more or less known in the literature; we mention, for the sake of the reader with experience in the Deift–Zhou method, that
  1. 1.
  2. 2.
We instead focus on the construction of the global (outer) parametrix which solves a model problem; the pragmatic reason is that this computation will produce a formula for the approximate location of the zeros of $\mathcal{T}_{n}(s)$ which can be actually tested numerically, and whose result is evident in Fig. 3. Some details on the construction of the local parametrices are contained in Section 5.2.3 and App. A.1.
5.2.1 Second transformation, model problem and its solution: the case $\mathrm{Re}\,(s)>0$
With reference to Fig. 8, left pane, we denote by $\Lambda_{\pm}^{(x,y)}$ the region bounded by $\Gamma_{x}^{y}$ and ±$\mathscr{L}_{\pm}^{(x,y)}$, for $x,y\in\{a_{+},a_{-},b,1\}$ and we call them lens regions (as opposed to their boundaries, which we refer to as the lens arcs).
Definition of $T_{0}$ for $\mathrm{Re}\,(s)>0$.
We refer to the left pane of Fig. 8 and then set
$T_{0}(z):=\left\{\begin{array}[]{lc}\displaystyle W(z)\begin{bmatrix}1&0\\ \frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix}&z\in\Lambda_{-}^{(a_{\pm},b)}\\[10.0pt] W(z)\begin{bmatrix}1&0\\ -\frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix}&z\in\Lambda_{+}^{(a_{\pm},b)}\end{array}\right..$ (5.44)
$T_{0}(z):=\left\{\begin{array}[]{lc}\displaystyle W(z)\begin{bmatrix}1&0\\ \frac{{\rm e}^{-n\varphi(z)-2i\pi{{\rho}}}}{\varkappa Q(z)}&1\end{bmatrix}&z\in\Lambda_{-}^{(b,1)}\\[10.0pt] W(z)\begin{bmatrix}1&0\\ -\frac{{\rm e}^{-n\varphi(z)}}{\varkappa Q(z)}&1\end{bmatrix}&z\in\Lambda_{+}^{(b,1)}\end{array}\right..$ (5.45)
The main arcs are as in Fig. 8, left pane, thus
$\Gamma_{m}=\Gamma_{b}^{1}\cup\Gamma_{a_{-}}^{b}\cup\Gamma_{b}^{a_{+}},\qquad\Gamma_{c}=\Gamma_{a_{+}}^{a_{-}}.$ (5.46)
Recall from Sec. 4.2 that we have defined
$\varphi(z;s):=\int_{a_{-}}^{z}\frac{2}{w^{2}}\sqrt{w^{2}+\frac{s^{2}}{4}+\frac{Aw^{2}}{w-1}}{\mathrm{d}}w,$ (5.47)
with the constant $A$ determined by the condition of all periods being purely imaginary and the integration path chosen in ${\mathbb{C}}\setminus\Gamma_{m}$. The effective potential satisfies the following boundary value relations
$\varphi(z_{+})+\varphi(z_{-})$ $=0$ $z\in\Gamma_{a_{-}}^{b}$ (5.48)
$\varphi(z_{+})+\varphi(z_{-})$ $=2\Omega_{1}$ $z\in\Gamma_{b}^{1}$ (5.49)
$\varphi(z_{+})+\varphi(z_{-})$ $=2\Omega_{2}$ $z\in\Gamma_{b}^{a_{+}}$ (5.50)
$\varphi(z_{+})-\varphi(z_{-})$ $=4i\pi$ $z\in\Gamma_{c},$ (5.51)
where the constants $\Omega_{1},\Omega_{2}$ are given by
$\Omega_{1}:=\int_{b}^{a_{+}}\varphi^{\prime}(z_{+}){\mathrm{d}}z\in i{\mathbb{R}}_{-},\qquad\Omega_{2}:=\int_{b}^{1}\varphi^{\prime}(z_{-}){\mathrm{d}}z\in i{\mathbb{R}}_{+}.$ (5.52)
From the Cauchy’s residue theorem one also has
$2\int_{a_{-}}^{b}\varphi^{\prime}(z_{-}){\mathrm{d}}z+2\int_{b}^{1}\varphi^{\prime}(z_{-}){\mathrm{d}}z+2\int_{b}^{a_{+}}\varphi^{\prime}(z_{-}){\mathrm{d}}z=4i\pi,$ (5.53)
with all the integrals in $i{\mathbb{R}}_{+}$.
A direct computation, using also the properties (5.48)–(5.51) shows that the matrix $T_{0}(z)$ satisfies the following RHP
Problem 5.7
The matrix $T_{0}(z)$ satisfies the conditions
$T_{0}(z_{+})=T_{0}(z_{-})J_{T_{0}}(z),\ \ \ \ z\in\gamma$ (5.54)
$T_{0}(z)=\mathbf{1}+\mathcal{O}\left(\frac{1}{z}\right),\ \ \ \ z\to\infty$ (5.55)
where
@@ 5.2.1 Second transformation, model problem and its solution: the case $\mathrm{Re}\,(s)>0$
$J_{T_{0}}(z)=\left\{\begin{array}[]{lc}J_{{}_{W}}(z)&z\in\Gamma_{c}=\Gamma_{a_{+}}^{a_{-}}\\[12.0pt] \begin{bmatrix}1&0\\ \frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix}&z\in\mathscr{L}_{+}^{(b,a_{+})}\cup\mathscr{L}_{-}^{(b,a_{+})}\cup\mathscr{L}_{-}^{(a_{-},b)}\cup\mathscr{L}_{+}^{(a_{-},b)}\\[15.0pt] \begin{bmatrix}1&0\\ \frac{{\rm e}^{-n\varphi(z)}}{\varkappa Q(z)}&1\end{bmatrix}&z\in\mathscr{L}_{+}^{(b,1)}\\[15.0pt] \begin{bmatrix}1&0\\ \frac{{\rm e}^{-n\varphi(z)}}{Q(z)}\left(1+\frac{1}{\varkappa}\right)&1\end{bmatrix}&z\in\mathscr{L}_{b}^{c_{+}}\\[15.0pt] \begin{bmatrix}1&0\\ \frac{{\rm e}^{-n\varphi(z)}}{Q(z)}\left(\frac{{\rm e}^{-2i\pi{{\rho}}}}{\varkappa}-1\right)&1\end{bmatrix}&z\in\mathscr{L}_{b}^{c_{-}}\\[15.0pt] \begin{bmatrix}1&0\\ \frac{{\rm e}^{-n\varphi(z)-2i\pi{{\rho}}}}{\varkappa Q(z)}&1\end{bmatrix}&z\in\mathscr{L}_{-}^{(b,1)}\\[15.0pt] \begin{bmatrix}0&\varkappa Q(z_{+}){\rm e}^{n\Omega_{1}}\\ -\frac{{\rm e}^{-n\Omega_{1}}}{\varkappa Q(z_{+})}&0\end{bmatrix}&z\in\Gamma_{b}^{1}\\[15.0pt] \begin{bmatrix}0&Q(z){\rm e}^{n\Omega_{2}}\\ -\frac{{\rm e}^{-n\Omega_{2}}}{Q(z)}&0\end{bmatrix}&z\in\Gamma_{b}^{a_{+}}\\[15.0pt] \begin{bmatrix}0&Q(z)\\ -\frac{1}{Q(z)}&0\end{bmatrix}&z\in\Gamma_{a_{-}}^{b}.\end{array}\right.$ (5.56)
Model problem.
The model problem that we have to solve is the following one
Problem 5.8

### v3

@@ 5.1.1 Conclusion of the steepest descent analysis
${\bf G}(z):=\begin{cases}M_{{}_{Q}}(z),&z\not\in\mathbb{D}_{+}\cup\mathbb{D}_{-},\\ \displaystyle M_{{}_{Q}}(z)Q(z)^{\frac{\sigma_{3}}{2}}\begin{bmatrix}1&-1\\ {\rm i}&{\rm i}\\ \end{bmatrix}\frac{\displaystyle\zeta_{\pm}(z)^{-\sigma_{3}/4}}{4\sqrt{\pi}}\mathcal{P}_{\pm}(z),&z\in\mathbb{D}_{\pm}.\end{cases}$
The matrix ${\bf G}(z)$ has the same jumps as $T_{0}(z)$ within $\mathbb{D}_{\pm}$ and on $\Gamma_{m}$. It has additional jump discontinuities across the boundaries $\partial\mathbb{D}_{\pm}$ of the form ${\bf G}(z_{+})={\bf G}(z_{-})\bigl(\mathbf{1}+\mathcal{O}\bigl(n^{-1}\bigr)\bigr)$, $z\in\partial\mathbb{D}_{\pm}$. Consequently, the matrix
@@
has only jump discontinuities on $\partial\mathbb{D}_{+}\cup\partial\mathbb{D}_{-}\cup\mathscr{L}_{+}\cup\mathscr{L}_{-}$ with all jump matrices that can be verified to be of the form $\mathbf{1}+\mathcal{O}\bigl(n^{-1}\bigr)$, see Figure 7.
It is then a standard result referred to in the literature as a “small norm theorem” [25, 26] that the solution of this last RHP $\mathscr{E}(z)$ exists and is uniformly close to the identity matrix. Reversing the order of the various transformations, this proves that $\forall s\in{\mathbb{C}}\setminus{\rm EoT}$, the solution $Y_{n}(z;s)$ of Riemann–Hilbert Problem 5.2 exists (at least for $n$ large enough).
More precisely, we can state as a result of the above discussion the following.
Theorem 5.8.
For any closed set $\mathcal{K}$ in the complement of the closure of $\rm EoT$}, there is $N_{0}\in{\mathbb{N}}$ such that the solution of Riemann–Hilbert Problem 5.2, and hence no zeros of $\mathcal{T}_{n}(s)$ (3.1) lie in $\mathcal{K}$ for $n\geq N_{0}$.
We only give a sketch of the proof, which is a relatively standard argument. We summarize the main steps here for the reader with some experience in the Deift–Zhou type of analysis.
As discussed previously in this section, the solvability of Riemann–Hilbert Problem 5.2 relies on the solvability of the RHP for the error matrix $\mathscr{E}(z)$ (5.8), whose jump matrices, $J_{\mathscr{E}}(z)$, can be bound uniformly, along the contours of the discontinuities, by
@@
The constant $C$ depends on the radii of the local disks $\mathbb{D}_{\pm}$, which is only limited by the distance of the branch points $\pm\frac{{\rm i}s}{2}$ from $z=1$, and it is bounded away from zero uniformly as $s\in\mathcal{K}$. The other source of potential loss of control over the error term is the distance of the main and complementary arcs $\Gamma_{m}$, $\Gamma_{c}$ from $z=1$. Indeed, if either arc contains $z=1$, the asymptotic analysis would require the construction of a special local parametrix and the small norm theorem cannot be used in the same way. However, as long as $s\in\mathcal{K}$ also the distance of $\Gamma_{m}\cup\Gamma_{c}$} from $z=1$ is bounded away from zero, and hence we can open the lenses as normal. Along the rims of the lenses, $\mathscr{L}_{\pm}$, and along $\Gamma_{c}$, the jump matrices of $\mathscr{E}$ tend exponentially to the identity, so that the main source of error is coming from the boundaries of $\mathbb{D}_{\pm}$.
Thus, in conclusion, as long as both branch points $\pm\frac{{\rm i}s}{2}$ and $\Gamma_{c}\cup\Gamma_{m}$ remains at finite positive distance from $z=1$ the constant $C$ in (5.9) can be chosen uniformly with respect to $s\in\mathcal{K}$ so that for $n$ sufficiently large the small norm theorem comes into effect, and uniformly so with respect to $s$.
@@
5.2 Asymptotic analysis inside $\boldsymbol{{\rm EoT}}$
Figure 8: The contours $\Gamma_{m}$ (red) and $\Gamma_{c}$ (green) for two antipodal values $s_{1}=-s_{2}$ of $s$ with $\operatorname{Re}(s_{1})>0$ (left) and $\operatorname{Re}(s_{2})<0$ (right). Specifically, here $s_{1}=0.48+0.48{\rm i}=-s_{2}$. The dashed line indicates the branch cut of the function $Q(z)$, which runs partially along $\Gamma_{b}^{1}$. The components of $\Gamma_{c}$ (green arcs) can be chosen freely within the shaded region, which indicates the region where $\operatorname{Re}\varphi<0$.
With reference to Figure 8, we have already observed that the level sets of $\operatorname{Re}\varphi(z;s)$ depend only on $s^{2}$ and hence are invariant under the map $s\mapsto-s$. The same is not true for the branch cuts $\Gamma_{m}$ and the regions of positivity of $\operatorname{Re}\varphi$. The decision of where to place the branch cuts is forced by the “sign distribution” (i.e., in which regions we have $\operatorname{Re}\varphi>0$). This is mandated by the behaviour near infinity $\operatorname{Re}\varphi=2\ln|z|+\mathcal{O}(1)$ for the outer region, and by the behaviour near the origin $\operatorname{Re}\varphi=\operatorname{Re}\frac{s}{z}+\mathcal{O}(1)$ for the two bounded regions. Before we fork the analysis according to the sign of $\operatorname{Re}(s)$, we make some common preparation.
As shown in Figure 8, one of the arcs of $\Gamma_{m}$ consists of an arc joining $z=1,b$, denoted $\Gamma_{b}^{1}$ (the orientation of the arc is suggested by the index placements). Thus, we need to preemptively modify the initial Riemann–Hilbert Problem 2.4 so that the contour $\gamma$ is “retracted” in part along $\Gamma_{b}^{1}$. We leave the details to the reader and we report the resulting RHP, which is now the starting point of the following analysis.
Remark 5.9.
Since the analysis is the complete parallel to that in Section 5.1, we recycle the symbols $M$, $W$, $T_{0}$ for the solutions of corresponding Riemann–Hilbert problems because, although they are different functions, they fulfill the same role. Since moreover we have to repeat the analysis for $\operatorname{Re}(s)>0$ and $\operatorname{Re}(s)<0$, we will recycle the symbols within each separate case.
Problem 5.10.
Let $\gamma=\Gamma_{b}^{1}\cup\gamma_{b}$ consist of the arc $\Gamma_{b}^{1}$ and a loop, $\gamma_{b}$, starting and ending at z=b{z=b} and containing the origin. Let $Q(z)$ in (5.1) be defined with the branch cut extending from $z=0$ to $z=b$ and then along $\Gamma_{b}^{1}$, as shown in Figure 8 by the dashed line. Given $n\in\mathbb{N}$, find a $2\times 2$} matrix-valued function $Y(z)=Y_{n}(z)$ such that
  1. $(1)$
@@ 2. $(2)$
    $Y(z_{+})=Y(z_{-})\begin{bmatrix}1&Q(z){\rm e}^{\frac{ns}{z}}\\ 0&1\end{bmatrix},\qquad\forall z\in\gamma_{b},$
    $Y(z_{+})=Y(z_{-})\begin{bmatrix}1&\varkappa Q(z_{+}){\rm e}^{\frac{ns}{z}}\\ 0&1\end{bmatrix},\qquad\forall z\in\Gamma_{b}^{1},\qquad\text{where}$
    $\varkappa:={\rm e}^{-2{\rm i}\pi{{\rho}}}-1.$ (5.10)
  3. $(3)$
    As $z\to\infty$, the matrix $Y_{n}(z)$ admits an asymptotic expansion of the form (2.6).
The first transformation is the same as in (5.2) and leads to the RHP for $W$ as follows.
Problem 5.11.
$W(z_{+})=W(z_{-})J_{W}(z),\qquad z\in\gamma,\qquad W(z)=\mathbf{1}+\mathcal{O}\left(\frac{1}{z}\right),\qquad z\to\infty,$
@@ 3. $(3)$
$J_{W}(z)=\begin{bmatrix}{\rm e}^{-\frac{n}{2}\left(\varphi(z_{+})-\varphi(z_{-})\right)}&Q(z){\rm e}^{\frac{n}{2}\left(\varphi(z_{+})+\varphi(z_{-})\right)}\\ 0&{\rm e}^{\frac{n}{2}\left(\varphi(z_{+})-\varphi(z_{-})\right)}\end{bmatrix},\qquad z\in\Gamma\setminus\Gamma_{b}^{1},$
$J_{W}(z)=\begin{bmatrix}{\rm e}^{-\frac{n}{2}\left(\varphi(z_{+})-\varphi(z_{-})\right)}&\varkappa Q(z_{+}){\rm e}^{\frac{n}{2}\left(\varphi(z_{+})+\varphi(z_{-})\right)}\\ 0&{\rm e}^{\frac{n}{2}\left(\varphi(z_{+})-\varphi(z_{-})\right)}\end{bmatrix},\qquad z\in\Gamma_{b}^{1}.$
The second transformation is also similar to (5.3). However, from this point onwards, the details of the transformation depend on the case $\operatorname{Re}(s)>0$ or $\operatorname{Re}(s)<0$ and thus are given in Sections 5.2.1 and 5.2.2, respectively.
We will not carry out completely the error analysis, which would require the construction of appropriate local parametrices near the points $z=b,a_{\pm},1$. These are more or less known in the literature. We mention, for the sake of the reader with experience in the Deift–Zhou method, that
  1. (1)
  2. (2)
We instead focus on the construction of the global (outer) parametrix which solves a model problem. The pragmatic reason is that this computation will produce a formula for the approximate location of the zeros of $\mathcal{T}_{n}(s)$ which can be actually tested numerically, and whose result is evident in Figure 3. Some details on the construction of the local parametrices are contained in Section 5.2.3 and Appendix A.1.
5.2.1 Second transformation, model problem and its solution: The case $\boldsymbol{\operatorname{Re}(s)>0}$
With reference to Figure 8, left pane, we denote by $\Lambda_{\pm}^{(x,y)}$ the region bounded by $\Gamma_{x}^{y}$ and $\mathscr{L}_{\pm}^{(x,y)}$, for $x,y\in\{a_{+},a_{-},b,1\}$ and we call them lens regions (as opposed to their boundaries, which we refer to as the lens arcs).
Definition of $\boldsymbol{T_{0}}$ for $\boldsymbol{\operatorname{Re}(s)>0}$. We refer to the left pane of Figure 8 and then set
$T_{0}(z):=\begin{cases}\displaystyle W(z)\begin{bmatrix}1&0\\ \displaystyle\frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix},&z\in\Lambda_{-}^{(a_{\pm},b)},\vskip 2.84526pt\\ W(z)\begin{bmatrix}1&0\\ \displaystyle-\frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix},&z\in\Lambda_{+}^{(a_{\pm},b)}.\end{cases}$
$T_{0}(z):=\begin{cases}\displaystyle W(z)\begin{bmatrix}1&0\\ \displaystyle\frac{{\rm e}^{-n\varphi(z)-2{\rm i}\pi{{\rho}}}}{\varkappa Q(z)}&1\end{bmatrix},&z\in\Lambda_{-}^{(b,1)},\vskip 2.84526pt\\ W(z)\begin{bmatrix}1&0\\ \displaystyle-\frac{{\rm e}^{-n\varphi(z)}}{\varkappa Q(z)}&1\end{bmatrix},&z\in\Lambda_{+}^{(b,1)}.\end{cases}$ (5.11)
The main arcs are as in Figure 8, left pane, thus
$\Gamma_{m}=\Gamma_{b}^{1}\cup\Gamma_{a_{-}}^{b}\cup\Gamma_{b}^{a_{+}},\qquad\Gamma_{c}=\Gamma_{a_{+}}^{a_{-}}.$
Recall from Section 4.2 that we have defined
$\varphi(z;s):=\int_{a_{-}}^{z}\frac{2}{w^{2}}\sqrt{w^{2}+\frac{s^{2}}{4}+\frac{Aw^{2}}{w-1}}{\mathrm{d}}w,$
with the constant $A$ determined by the condition of all periods being purely imaginary and the integration path chosen in ${\mathbb{C}}\setminus\Gamma_{m}$. The effective potential satisfies the following boundary value relations:
$\varphi(z_{+})+\varphi(z_{-})=0,\qquad z\in\Gamma_{a_{-}}^{b},$ (5.12)
$\varphi(z_{+})+\varphi(z_{-})=2\Omega_{1},\qquad z\in\Gamma_{b}^{1},$ (5.13)
$\varphi(z_{+})+\varphi(z_{-})=2\Omega_{2},\qquad z\in\Gamma_{b}^{a_{+}},$ (5.14)
$\varphi(z_{+})-\varphi(z_{-})=4{\rm i}\pi,\qquad z\in\Gamma_{c},$ (5.15)
where the constants $\Omega_{1}$, $\Omega_{2}$ are given by
$\Omega_{1}:=\int_{b}^{a_{+}}\varphi^{\prime}(z_{+}){\mathrm{d}}z\in{\rm i}{\mathbb{R}}_{-},\qquad\Omega_{2}:=\int_{b}^{1}\varphi^{\prime}(z_{-}){\mathrm{d}}z\in{\rm i}{\mathbb{R}}_{+}.$ (5.16)
From Cauchy’s residue theorem, one also has
$2\int_{a_{-}}^{b}\varphi^{\prime}(z_{-}){\mathrm{d}}z+2\int_{b}^{1}\varphi^{\prime}(z_{-}){\mathrm{d}}z+2\int_{b}^{a_{+}}\varphi^{\prime}(z_{-}){\mathrm{d}}z=4{\rm i}\pi,$
with all the integrals in $\rm i$$\mathbb{R}$$_{+}$.
A direct computation, using also the properties (5.12)–(5.15) shows that the matrix $T_{0}(z)$ satisfies the following RHP.
Problem 5.12.
The matrix $T_{0}(z)$ satisfies the conditions
$T_{0}(z_{+})=T_{0}(z_{-})J_{T_{0}}(z),\qquad z\in\gamma,\qquad T_{0}(z)=\mathbf{1}+\mathcal{O}\left(\frac{1}{z}\right),\qquad z\to\infty,$
@@ 5.2.1 Second transformation, model problem and its solution: The case $\boldsymbol{\operatorname{Re}(s)>0}$
$J_{T_{0}}(z)=\begin{cases}J_{{}_{W}}(z),&z\in\Gamma_{c}=\Gamma_{a_{+}}^{a_{-}},\\ \begin{bmatrix}1&0\\ \displaystyle\frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix},&z\in\mathscr{L}_{+}^{(b,a_{+})}\cup\mathscr{L}_{-}^{(b,a_{+})}\cup\mathscr{L}_{-}^{(a_{-},b)}\cup\mathscr{L}_{+}^{(a_{-},b)},\vskip 2.84526pt\\ \begin{bmatrix}1&0\\ \displaystyle\frac{{\rm e}^{-n\varphi(z)}}{\varkappa Q(z)}&1\end{bmatrix},&z\in\mathscr{L}_{+}^{(b,1)},\vskip 2.84526pt\\ \begin{bmatrix}1&0\\ \displaystyle\frac{{\rm e}^{-n\varphi(z)}}{Q(z)}\left(1+\frac{1}{\varkappa}\right)&1\end{bmatrix},&z\in\mathscr{L}_{b}^{c_{+}},\vskip 2.84526pt\\ \begin{bmatrix}1&0\\ \displaystyle\frac{{\rm e}^{-n\varphi(z)}}{Q(z)}\left(\frac{{\rm e}^{-2{\rm i}\pi{{\rho}}}}{\varkappa}-1\right)&1\end{bmatrix},&z\in\mathscr{L}_{b}^{c_{-}},\vskip 2.84526pt\\ \begin{bmatrix}1&0\\ \displaystyle\frac{{\rm e}^{-n\varphi(z)-2{\rm i}\pi{{\rho}}}}{\varkappa Q(z)}&1\end{bmatrix},&z\in\mathscr{L}_{-}^{(b,1)},\vskip 2.84526pt\\ \begin{bmatrix}0&\varkappa Q(z_{+}){\rm e}^{n\Omega_{1}}\\ \displaystyle-\frac{{\rm e}^{-n\Omega_{1}}}{\varkappa Q(z_{+})}&0\end{bmatrix},&z\in\Gamma_{b}^{1},\vskip 2.84526pt\\ \begin{bmatrix}0&Q(z){\rm e}^{n\Omega_{2}}\\ \displaystyle-\frac{{\rm e}^{-n\Omega_{2}}}{Q(z)}&0\end{bmatrix},&z\in\Gamma_{b}^{a_{+}},\vskip 2.84526pt\\ \begin{bmatrix}0&Q(z)\\ \displaystyle-\frac{1}{Q(z)}&0\end{bmatrix},&z\in\Gamma_{a_{-}}^{b}.\end{cases}$ (5.17)
Model problem. The model problem that we have to solve is the following one.
Problem 5.13.